OpenStudy (anonymous):

NEED HELP!! A soccer team wins 65% of its matches, and 15% of its matches end in a draw. If the team is scheduled to play 20 matches, about how many matches is it expected to lose? 13 4 8 1

OpenStudy (anonymous):

@shanemc1010 please help? I will medal

OpenStudy (anonymous):

and fan if you help me with other questions too

OpenStudy (anonymous):

okay let me think

OpenStudy (anonymous):

100-65=percentage lost

OpenStudy (anonymous):

100-65=35 so what do I have to do now

OpenStudy (anonymous):

i got 100−65=?35 35=35

OpenStudy (anonymous):

yeah me too and it's so weird cause that's not one of the options?

OpenStudy (anonymous):

what are your options?

OpenStudy (anonymous):

13 4 8 1

OpenStudy (anonymous):

k hold on let me go over it

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i would go with 1 or 4

OpenStudy (anonymous):

is there one answer that you're very sure on?

OpenStudy (anonymous):

someone pleaaaasee help

OpenStudy (anonymous):

its 4 dudes

OpenStudy (anonymous):

how'd you do it?

OpenStudy (anonymous):

and yes ur correct

OpenStudy (anonymous):

sorry about that i was at working

OpenStudy (anonymous):

im back do you have any more questions

OpenStudy (anonymous):

by the way when i put 35 that is the percent they lost

OpenStudy (anonymous):

not the exact number of games

OpenStudy (anonymous):

ohhh okay and not for now but in a little bit I'll have more, thanks :)